Q21:In a survey of 444 HIV-positive smokers, 190 reported that they had used a nicotine patch to try to quit smoking. Can you conclude that less than half of HIV-positive smokers have used a nicotine patch? Use the a=0.10 level of significance and the P-value method with the TI-84 Plus calculator.

A) H0:

H1:

right tailed, left tailed, or two tailed?

B)Compute the value of the test statistic.

C) Using a=0.10, can you conclude that less than half of HIV-positive smokers have used a nicotine patch?

D) State a conclusion
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

a.

p = x/n = 190/444 = .4279

Ho : po >= 0.5

Ha : po < 0.5

This is a left tailed test

b. test statistic = (p^ - p0)/sqrt(p0*p0'/n) = (.4279-.50)/sqrt(.5*.5/444) = -3.04

The p-value is .001

c. alpha = 0.10, p-value is less than alpha - reject Ho

d. We conclude that that less than half of HIV-positive smokers have used a nicotine patch